OpenStudy (anonymous):

-10 = -2 + 1/2 d solve for d

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-10+2=(1/2)d -8=(1/2)d -16=d d=-16

OpenStudy (anonymous):

are you shure this time???

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

*sure

OpenStudy (anonymous):

ji........lock it you can confirm also by substituting the value of d in question.....

OpenStudy (anonymous):

any other question?